Let Z~N (0,1).Find a constant c for which P(-c≤ Z ≤ c) = 0.8664


1
Expert's answer
2021-07-23T12:01:01-0400

By condition, the mathematical expectation is "a=0" and the standard deviation is "\sigma = 1" .

Then

"P( - c \le Z \le c) = 0.8664 \Rightarrow \Phi \left( {\frac{{c - a}}{\sigma }} \right) - \Phi \left( {\frac{{ - c - a}}{\sigma }} \right) = 0.8664 \Rightarrow \Phi \left( c \right) - \Phi \left( { - c} \right) = 0.8664 \Rightarrow 2\Phi \left( c \right) = 0.8664 \Rightarrow \Phi \left( c \right) = 0.4332 \Rightarrow c = 1.5"

Answer: "c=1.5"